OpenStudy (anonymous):

Antarctica comprises about 10.5% of the total land area on Earth. Antarctica has a surface area of about 5,400,000 square miles. What is the total land area on Earth?

OpenStudy (anonymous):

5,400,000/10.5 = x/100

OpenStudy (anonymous):

Antartica comprises about 10.5% of the total land area on earth. Antartica has a surface area of about 5,400,000 square miles. What is the total land area on Earth?

OpenStudy (anonymous):

5,400,000/10.5 = x/100